Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,000 units × $20 per unit) $ 260,000
Variable expenses 130,000
Contribution margin 130,000
Fixed expenses 145,000
Net operating loss $ (15,000 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,700 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $87,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $39,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.70 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,800?

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $58,000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

b. Assume that the company expects to sell 20,100 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,100)?
Answer #1

1. Sales price per unit = $ 20, for 13,000 units , so for Contribution per unit = $ 10 ( 130,000/13000)

Company CM ratio = Contribution margin / sales

CM Ratio = $ 130,000 / 260,000

CM Ratio = 50%.

Break even point in unit sales = Fixed cost / contribution margin per unit.

Break even point in units sales = $ 145,000 / 10

Break even point in unit sales = 14,500 units.

Break even point in dollar sales = Fixed cost / contribution margin%. = $ 145,000 / 50%

Break even point in dollar sales = $ 290,000.

2.

Sales increase $ 87,000 so revised sales = $ 347,000. And revised fixed cost = $ 151700

Thus contribution margin for this sales level = $ 347,000 × 50% = $ 173,500.

Operating profit = Contribution margin - Fixed cost

Operating profit = $ 173,500 - $ 151,700 = $ 21,800.

Thus Increase in Net profit = Existing Profit - Revised profit

Increase in net operating income = - 15,000 - 21800 = $ 36,800.

3.

Selling price = $ 20 - 10% = $ 18 , sales units = 2 × 13000 = 26000 units and fixed cost = $ 184,000.

Revised sales = Units × sales price = 26000 × $ 18 = $ 468,000.

Contribution margin = $ 468,000 × 50% = $ 234,000 ( assumed CM ratio is not changed ).

Operating income = Contribution margin - Fixed cost = $ 234,000 - $ 184,000

Net operating income = $ 50,000.

Increase in net operating income = - 15,000 - 50000 = $ 65,000

4.

Packaging variable cost Increases by $ 0.7 per unit , so variable cost per unit = $ 10.70 ( $ 10 + 0.7)

Thus contribution margin per unit = $ 20 - 10.7 = $ 9.30

Units for Target profit = ( target profit + fixed cost ) / contribution margin per unit.

Units for Target profit = ( $ 4800 + $ 145,000) / 9.30

Units for Target profit = 16107.53 i.e 16108 units.

Thus units would have to be sold each month to attain a target profit of $4,800 = 16108 units ( rounded off)

As per HomeworkLib's policy first 4 questions to be answered if no questions are asked specifically...the rest of the question follows the same pattern of solutions.

Feel free to comment doubts if any..